OpenStudy (anonymous):

If the original coordinate axes are rotated 30° to obtain the x' and y' axes, what is the value of x in terms of x' and y'?

OpenStudy (anonymous):

\[x=\sqrt{\frac{ 3 }{ 2}}x'-\frac{ 1 }{ 2}y'\]

OpenStudy (anonymous):

\[x=\sqrt{\frac{ 3 }{ 2}}x'+\frac{ 1 }{ 2}y'\]

OpenStudy (anonymous):

x=\\frac{ 1 }{ 2}}x'-\[x=\frac{ 1 }{ 2}x'-\sqrt{\frac{ 3 }{ 2}}y'\]\frac{ 1 }{ 2}y'

OpenStudy (anonymous):

\[x=\frac{ 1}{ 2}x'+\sqrt \frac{ 3 }{ 2}y'\]

OpenStudy (anonymous):

a b c d

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

@dan815

OpenStudy (michele_laino):

here we can compute the (2x2) matrix N such that the subsequent equation holds: \[\left( {\begin{array}{*{20}{c}} {x'} \\ {y'} \end{array}} \right) = N\left( {\begin{array}{*{20}{c}} x \\ y \end{array}} \right)\]